The mean lives of radio active substances are 1620 years and 405 years for a particle and b particle respectively. Find out the time during three fourth of a sample will decay if it is decaying by both a and b emission simultaneously.

Grade:12

1 Answers

Vikas TU
7 years ago
Total decal constant that is,
k = k1k2/(k1+k2) =>after putting the values k1 = 1620 and 405
net k = 1/324 year.
N/4 would left after disintegrating three-fourth and hence,
N/4 = e^-(kt)
taking log both sides we get,
t = (2/k)ln(2) = > 449 years. approx.